silver ,copper ,lead and zinc are melted together in a ratio of 2:6:7:8 and formed into a mass . one kg of zinc cost Rs 15, a kg of lead 20% more than that of zinc,a kg of silver 20 times that of lead . if the quantity melted was sold at Rs.60 per kg,find the price of copper in rupees per kilogram.

Answer by ankor@dixie-net.com(22740)

silver ,copper ,lead and zinc are melted together in a ratio of 2:6:7:8 and formed into a mass .
one kg of zinc cost Rs 15,
a kg of lead 20% more than that of zinc,
a kg of silver 20 times that of lead .
if the quantity melted was sold at Rs.60 per kg, find the price of copper in rupees per kilogram.
:
let's use the sum of the ratio values as the resulting total 23 kg
then
2 km, amt of silver
6 kg, amt of copper
7 kg, amt of lead
8 kg, amt of zinc
:
let c = cost of copper per kg
Given: Zinc = $15 per kg
lead: 1.2(15) = $18 per kg
silver; 20(18) = $360 per kg
:
The mixture equation
360(2) + 6c + 18(7) + 15(8) = 60(23)
720 + 6c + 126 + 120 = 1380
6c + 966 = 1380
6c = 1380 - 966
6c = 414
c = 414/6
c = $69 is the cost per kg for copper